OpenStudy (anonymous):

Solve for x, the equation 12x^3+8x^2-x-1=0. (All solutions are rational and between +-1.)

OpenStudy (anonymous):

@jim_thompson5910 @Loser66

OpenStudy (anonymous):

(2x + 1 ) (6x^2 + x -1 ) = 0

OpenStudy (anonymous):

then ( 6x^2 + x -1 ) = 0 well equal ( 3x +1) (2x -1 ) =0 so x = -0.5 , -1/3 , =1/2